A car traveling 88 km/hr is 110 m behind a truck traveling 75 km/hr. How long will it take the car to reach the truck?
Answer by oberobic(2304)

I assume the car is 110 km behind the truck.
The difference in their speeds is 88-75 = 13 km/hr.
The car needs to make up 110 km.
t = d/r
t = 110/13
t = 8.46 hr for the car to catch up
8.46 hr = about 8 hr 28 min